A typical incandescent bulb has an efficiency of about 10%. What current is needed for a bulb with a resistance of (1.5×10^2) Ω  to output power of 6.0 W?


1
Expert's answer
2022-03-14T13:06:22-0400

Solution;

"P=I^2R"

"6=I^2(1.5\u00d710^2)"

"I^2=0.4"

"I=0.6324A"